Demand is given by Q(p) = 530-2p. What is the price elasticity of demand when p=100? p=200? Please show work
Answer #1

Answer

Price elasticity of demand =(dQ/dP)*(P/Q)

dQ/dP=-2 .......... the first differentiation of the demand curve concerning P by power rule.

P=100

Q=530-2*100=330

elasticty =(-2)*(100/330)=-0.606060606=-0.61

the demand is inelastic as the elastcity is above -1.

--------

P=200

Q=530-2*200=130

Elastcity =(-2)*(200/130)=-3.07692308=-3.08

the demand is elastic as the elastcity is below -1.